Julius sold five times as many compters as Sam sold last year.In total, they sold 78 computers.How many computers did julius sell?
Answer by checkley77(12844)

x+5x=78
6x=78
x=78/6
x=13 computers sold by Sam.
5*13=65 computers sold by Julius.